The correct answer to 2 ×107 × 8 ×10-12 in standard form is:

A) 1.6 x 10-4

B) 16 x 1019

C) 1.6 x 1019

D) 16 x 10-5

1
Expert's answer
2022-06-02T10:26:34-0400
"2 ×10^7 × 8 ×10^{-12}=16\times 10^{-5}=1.6\times 10^{-4}"

Answer: A) 1.6 x 10-4